OpenStudy (anonymous):

Simplify the expression by combining like terms. x - 24y + x + 24y is it 48y?

OpenStudy (campbell_st):

look at the terms in x x + x = 2x -24y + 24y = 0 then x + x -24y + 24y = 2x

OpenStudy (kymber):

 Remember that the signs BEFORE the number goes with it!